# On the quantitative lower bounds for solutions to the Boltzmann equation in nonconvex domains

### v3

@@ 6. Lower bound for the non-cutoff case
(6.22) $\begin{split}|Q_{\epsilon}^{1}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)|\leq C_{Q,2}m_{b_{NCO}}c_{\Phi}E^{\prime}_{f}W_{f},\end{split}$
@@
(6.23) $\begin{split}&f(t,x,v)\\ &\geq\int_{\tau}^{t}\exp\left(-(t-s)(C_{L,1}m_{b_{CO}}+C_{L,2}m_{b_{NCO}})\langle|v^{\prime}|+2r_{k}(\Delta_{2})\rangle^{\gamma^{+}}\right)\\ &[C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}}-C_{Q,2}m_{b_{NCO}}c_{\Phi}E^{\prime}_{f}W_{f}]ds\\ &\geq\int_{\tau}^{t}\exp\left(-(t-s)(C_{L,1}m_{b_{CO}}+C_{L,2}m_{b_{NCO}})\langle|v^{\prime}|+2r_{k}(\Delta_{2})\rangle^{\gamma^{+}}\right)\\ &C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}}ds-tC_{Q,2}m_{b_{NCO}}c_{\Phi}E^{\prime}_{f}W_{f},\end{split}$
@@
Then, we notice that
(6.24) $l_{b^{CO}_{\epsilon}}\geq l_{b},$
and that
(6.25) $\begin{split}n_{b^{CO}_{\epsilon}}:=\int_{\mathbb{S}^{2}}b^{CO}_{\epsilon}(\cos{\theta})d\sigma\sim\begin{cases}\frac{b_{0}}{\nu}\epsilon^{-\nu},\,\nu\in(0,2),\\ b_{0}|\log{\epsilon}|,\,\nu=0,\end{cases}\end{split}$
(6.26) $\begin{split}m_{b^{NCO}_{\epsilon}}:=\int_{\mathbb{S}^{2}}b^{NCO}_{\epsilon}(\cos{\theta})(1-\cos{\theta})d\sigma\sim\begin{cases}\frac{b_{0}}{2-\nu}\epsilon^{2-\nu},\,\nu\in(0,2),\\ \frac{b_{0}}{2}\epsilon^{2},\,\nu=0.\end{cases}\end{split}$
@@
(6.27) $tC_{2}m_{b^{NCO}_{\epsilon}}c_{\Phi}E^{\prime}_{f}W_{f}\leq C_{2}m_{b^{NCO}_{\epsilon}}c_{\Phi}E^{\prime}_{f}W_{f}\leq\frac{1}{2}C_{1}\alpha^{2}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)l_{b}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}},$
@@
(6.28) $\begin{split}&f(t,x,v)\\ &\geq\frac{1}{2}\int_{\max\left\{\tau,t-\frac{\Delta_{2}}{2^{k+1}(2r_{k}(\Delta_{2})+|v^{\prime}|)}\right\}}^{t}\exp\left(-(t-s)C_{L}\langle|v^{\prime}|+2r_{k}(\Delta_{2})\rangle^{\gamma^{+}}\right)\\ &\alpha^{2}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)C_{Q}(r_{k}(\Delta_{2}))^{3+\gamma}\xi^{\frac{1}{2}}ds,\end{split}$
for $v\in B(v^{\prime},r_{k}(\Delta_{2})\sqrt{2}(1-\xi))$.
We conclude the proof by setting $\xi=\frac{1}{4}$ and notice as in the cutoff case that $B(v^{\prime},r_{k+1}(\Delta_{2})))\subset B(0,|v^{\prime}|+2r_{k}(\Delta_{2}))$.
@@
Given $\Omega\subset\mathbb{R}^{3}$ satisfying Assumption A, a kernel $B$ satisfying Assumption B with $\nu\geq 0$, $\alpha\in[0,1]$, $\delta$, $d>0$ in Lemma 2.1, we consider a continuous mild solution $f(t,x,v)$ of (1.1)–(1.3) and $\tau\in(0,\Delta^{0})$, where $\Delta^{0}>0$ is given in Proposition 3.1. Then, there exist $b(\tau)$, $\delta_{T}(\tau)>0$, which depend on $\gamma,\ \nu,\ b_{0},\ n_{b}$, $m_{b}$ $C_{\Phi}$ and $E_{f}$, $W_{f}$, $E^{\prime}_{f}$, and MM(and $L_{f,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that for $t\in[\tau,\tau+\delta_{T}(\tau)]$,
(6.29) $\forall x\in\partial\Omega,\ \int_{v_{*}\cdot n(x)>0}f(t,x,v_{*})(v_{*}\cdot n(x))dv_{*}>b(\tau).$
@@
We consider $\Omega\subset\mathbb{R}^{3}$ satisfying Assumption A, kernel $B$ satisfying Assumption B with $\nu\geq 0$, $\alpha\in[0,1]$. We consider a continuous mild solution $f(t,x,v)$ of (1.1)–(1.3), namely, satisfying (6.3) and (6.4), and $\tau\in(0,\Delta^{0})$, where $\Delta^{0}>0$ is given in Proposition 3.1, $\{\xi_{n}\}_{n=1}^{\infty}\in(0,1)$. Then, there exist $\{\Delta_{i}\}_{i=1}^{\infty}$ such that $\sum_{i=1}^{\infty}\Delta_{i}=1$ and that for any $n\in\mathbb{N}$, $t\in[(\sum_{i=1}^{n}\Delta_{i})\tau,\tau]$, $\forall(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$, we have
(6.30) $f(t,x,v)\geq a_{n}(\tau)\mathbf{1}_{B(0,r_{n})}(v)$
@@
(6.31) $r_{n+1}:=\sqrt{2}r_{n}(1-\xi_{n}),$
(6.32) $a_{n+1}(\tau):=\begin{cases}C_{Q}\Delta_{n+1}e^{-[C_{L}a_{n}^{2}(\tau)r^{3+\gamma-\widetilde{\gamma}}\xi_{n}^{\frac{1}{2}}]^{\frac{-\nu}{2-\nu}}(\sum_{k\geq n+1}\Delta_{k})r^{\gamma^{+}}_{n}}a^{2}_{n}(\tau)r_{n}^{\gamma+3}\xi_{n}^{\frac{5}{2}},\nu\in(0,2),\\ C_{Q}\Delta_{n+1}e^{-C\log[C_{L}a_{n}^{2}(\tau)r^{3+\gamma-\widetilde{\gamma}}\xi_{n}^{\frac{1}{2}}](\sum_{k\geq n+1}\Delta_{k})r^{\gamma^{+}}_{n}}a^{2}_{n}(\tau)r_{n}^{\gamma+3}\xi_{n}^{\frac{5}{2}},\nu=0.\end{cases}$
@@
7. Appendix: the characteristic line
In this section, we recall some useful properties about the bounce trajectory of $\Omega$ for the definition of the characteristic line $X_{s,t}(x,v),\ V_{s,t}(x,v)$. To begin with, we introduce some notations from Definition A.1 in [2].
Definition 7.1.
@@ 7. Appendix: the characteristic line
Thanks to Proposition A.2 in [2], there is no specular reflection trajectory that leads to any point in $\Omega_{stop}$. Given $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$, we define
$t_{min}(x,v):=\max\{t\geq 0\mid x-vs\in\overline{\Omega},\forall\,0\leq s\leq t\}.$
@@
(1) If there exists $t$ in $(0,t_{min}(x,v))$ such that $x-vt\in\partial\Omega$, then $(x-vt,v)\in\Omega_{rolling}$
(2) $t_{min}=0$ if and only if $(x,v)\in\Omega_{stop}\cup\Omega_{rebound}$.
(3) $(x-vt_{min}(x,v),v)\in\Omega_{stop}\cup\Omega_{rebound}$
@@
Proposition 7.2.
Given an open bounded domain in $\mathbb{R}^{3}$ with $C^{1}$ boundary and $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$, for any $t\geq 0$ the trajectory of (x,v) with specular reflection after time t has at most countable number of rebound and rolling. Moreover, the set with infinite rebound and rolling in finite time is zero measure in $\overline{\Omega}\times\mathbb{R}^{3}$.
Remark 12.
@@
Now, we shall introduce the characteristic line by defining:
(7.1) $(t_{0},x_{0},v_{0}):=(0,x,v),$
(7.2) $(t_{k+1},x_{k+1},v_{k+1}):=\begin{cases}(\infty,x_{k},v_{k}),\ (x_{k},v_{k})\in\Omega_{stop},\\ (t_{k}+t_{min}(x_{k},v_{k}),x_{k}-v_{k}t_{min}(x_{k},v_{k}),R(x_{k+1},v_{k})),\ (x_{k},v_{k})\notin\Omega_{stop},\end{cases}$
(7.3) $n(t,x,v):=\sup\{k\in\mathbb{N}:\ t_{k}(x,v)\leq t\}.$
@@
Next, we define the last rebound from backward trajectory of $(x,v)$:
(7.4) $(t_{fin},x_{fin},v_{fin}):=\begin{cases}(t,x_{n},v_{n}),&n(t,x,v)<\infty,\ \,t_{n(t,x,v)+1}=\infty,\\ (t_{n},x_{n},v_{n}),&n(t,x,v)<\infty,\ \,t_{n(t,x,v)+1}<\infty,\\ \lim\limits_{n\rightarrow\infty}(t_{n},x_{n},v_{n}),&n(t,x,v)=\infty.\end{cases}$
Finally, we can define the characteristic line for $0\leq s\leq t$:

### v4

@@ 6. Lower bound for the non-cutoff case
(6.25) $\begin{split}|Q_{\epsilon}^{1}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)|\leq C_{Q,2}m_{b^{NCO}_{\epsilon}}c_{\Phi}E^{\prime}_{f}W_{f},\end{split}$
@@
(6.26) $\begin{split}&f(t,x,v)\\ \geq&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636pt\left(C_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\frac{1}{2}-C_{Q,2}m_{b^{NCO}_{\epsilon}}c_{\Phi}E^{\prime}_{f}W_{f}\right)\,ds\\ \geq&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\frac{1}{2}\,ds\\ -&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon}}+C_{L,2}m_{b^{NCO}_{\epsilon}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,2}m_{b^{NCO}_{\epsilon}}c_{\Phi}E^{\prime}_{f}W_{f}\,ds,\end{split}$
@@
Then, we notice that when $\frac{\pi}{4}>\epsilon>0$, we ahve
(6.27) $l_{b^{CO}_{\epsilon}}:=\inf\limits_{\frac{1}{4}\pi\leq\theta\leq\frac{3}{4}\pi}b^{CO}_{\epsilon}(\cos{\theta})\geq l_{b},$
Thus, we deduce from (6.4) and (6.5) that
(6.28) $\begin{split}n_{b^{CO}_{\epsilon}}=\int_{\mathbb{S}^{2}}b^{CO}_{\epsilon}(\cos{\theta})\,d\sigma\sim\begin{cases}\frac{b_{0}}{\nu}\epsilon^{-\nu},\,\nu\in(0,2),\\ b_{0}|\log{\epsilon}|,\,\nu=0,\end{cases}\end{split}$
(6.29) $\begin{split}m_{b^{NCO}_{\epsilon}}=\int_{\mathbb{S}^{2}}b^{NCO}_{\epsilon}(\cos{\theta})(1-\cos{\theta})\,d\sigma\sim\begin{cases}\frac{b_{0}}{2-\nu}\epsilon^{2-\nu},\,\nu\in(0,2),\\ \frac{b_{0}}{2}\epsilon^{2},\,\nu=0.\end{cases}\end{split}$
@@
(6.30) $\begin{split}&\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon_{k}}}+C_{L,2}m_{b^{NCO}_{\epsilon_{k}}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,2}m_{b^{NCO}_{\epsilon_{k}}}c_{\Phi}E^{\prime}_{f}W_{f}\,ds\\ \leq&\frac{1}{4}\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon_{k}}}+C_{L,2}m_{b^{NCO}_{\epsilon_{k}}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon_{k}}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\,ds\end{split}$
@@
(6.31) $\begin{split}&f(t,x,v)\\ \geq&\frac{1}{4}\int_{t^{NC}_{k}(\tau,t,\Delta_{2},|v^{\prime}|)}^{t}\exp\left(-(t-s)(C_{L,1}n_{b^{CO}_{\epsilon_{k}}}+C_{L,2}m_{b^{NCO}_{\epsilon_{k}}})\left\langle|v^{\prime}|+2r_{k}(\Delta_{2})\right\rangle^{\gamma^{+}}\right)\\ &\hskip 14.22636ptC_{Q,1}\alpha^{NC}_{k}(\tau,s,\Delta_{2},A,|v^{\prime}|)^{2}l_{b^{CO}_{\epsilon_{k}}}c_{\Phi}(r_{k}(\Delta_{2}))^{3+\gamma}\,ds\\ =&\alpha^{NC}_{k+1}(\tau,t,\Delta_{2},A,|v^{\prime}|),\end{split}$
for $v\in B(v^{\prime},r_{k+1}(\Delta_{2}))$ and we conclude the proof by induction.
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the kernel $B$ satisfies Assumption B with $\nu\geq 0$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha\in[0,1]$. Fix $\tau\in(0,\Delta^{0})$, with $\Delta^{0}>0$ is given in Proposition 3.1. Then, there exist $b^{NC}(\tau)$, $\delta_{T}(\tau)>0$, which depend on $\gamma,\ \nu,\ b_{0},\ n_{b}$, $m_{b}$ $C_{\Phi}$ and $E_{f}$, $W_{f}$, $E^{\prime}_{f}$, and MM(and $L_{f,p}$, where $p>\frac{3}{3+\gamma}$, if $\gamma<0$) such that for $t\in[\tau,\tau+\delta_{T}(\tau)]$,
(6.32) $\forall x\in\partial\Omega,\ \int_{v_{*}\cdot n(x)>0}f(t,x,v_{*})(v_{*}\cdot n(x))\,dv_{*}>b^{NC}(\tau).$
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the kernel $B$ satisfies Assumption B with $\nu\geq 0$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha\in[0,1]$. Fix $\tau\in(0,\Delta^{0})$, where $\Delta^{0}>0$ is given in Proposition 3.1 and $\{\xi_{n}\}_{n=1}^{\infty}\in(0,1)$. Then, there exist $r^{NC}_{0}>0$, $\{\Delta_{i}\}_{i=1}^{\infty}$ such that $\sum_{i=1}^{\infty}\Delta_{i}=1$ and that for any $n\in\mathbb{N}$, $t\in[(\sum_{i=1}^{n}\Delta_{i})\tau,\tau]$, $\forall(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$, we have
(6.33) $f(t,x,v)\geq a_{n}^{NC}(\tau)\mathbf{1}_{B(0,r^{\prime\prime}_{n})}(v)$
@@
(6.34) $r^{NC}_{n+1}:=\sqrt{2}\,(1-\xi_{n})\,r^{NC}_{n},$
(6.35) $a_{n+1}^{NC}(\tau):=\begin{cases}C_{Q}\Delta_{n+1}e^{-\left[C_{L}\left(a^{NC}_{n}(\tau)\right)^{2}\left(r_{n}^{NC}\right)^{3+\gamma-\widetilde{\gamma}}\xi_{n}^{\frac{1}{2}}\right]^{\frac{-\nu}{2-\nu}}\left(\sum\limits_{k\geq n+1}\Delta_{k}\right)\left(r_{n+1}^{NC}\right)^{\gamma^{+}}}\\ \times\left(a^{NC}_{n}(\tau)\right)^{2}\left(r_{n}^{NC}\right)^{\gamma+3}\xi_{n}^{\frac{5}{2}},&\nu\in(0,2),\\ C_{Q}\Delta_{n+1}e^{-C\log\left[C_{L}\left(a^{NC}_{n}(\tau)\right)^{2}\left(r_{n}^{NC}\right)^{3+\gamma-\widetilde{\gamma}}\xi_{n}^{\frac{1}{2}}\right]\left(\sum\limits_{k\geq n+1}\Delta_{k}\right)\left(r_{n+1}^{NC}\right)^{\gamma^{+}}}\\ \times\left(a^{NC}_{n}(\tau)\right)^{2}\left(r_{n}^{NC}\right)^{\gamma+3}\xi_{n}^{\frac{5}{2}},&\nu=0.\end{cases}$
@@
A. Appendix: the characteristic line
In this section, we recall some useful properties about the bounce trajectory of $\Omega$ used in the definition of the characteristic line $X_{s,t}(x,v)$ and $V_{s,t}(x,v)$. To begin with, we introduce some notations from Definition A.1 in [2].
Definition A.1.
@@ A. Appendix: the characteristic line
Due to Proposition A.2 in [2], there is no specular reflection trajectory that leads to any point in $\Omega_{stop}$. Given $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$, we recall the definition of $t_{\partial}(x,v)$
$t_{\partial}(x,v)=\max\{t\geq 0\mid x-vs\in\overline{\Omega},\forall\,s\in[0,t]\}.$
@@
(1) If there exists $t\in(0,t_{\partial}(x,v))$ such that $x-vt\in\partial\Omega$, then $(x-vt,v)\in\Omega_{rolling}$.
(2) $t_{\partial}(x,v)=0$ if and only if $(x,v)\in\Omega_{stop}\cup\Omega_{rebound}$.
(3) $(x-vt_{\partial},v)\in\Omega_{stop}\cup\Omega_{rebound}$.
@@
Proposition A.2.
We consider an open bounded domain in $\mathbb{R}^{3}$ with $C^{1}$ boundary and $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$. Then, for any $t\geq 0$ the trajectory of $(x,v)$ with specular reflection after time $t$ has at most a countable number of rebound and rolling. Moreover, the set with infinite rebound and rolling in finite time is zero measure in $\overline{\Omega}\times\mathbb{R}^{3}$.
Remark A.1.
@@
Now, we introduce the characteristic line by defining:
(A.1) $(t_{0},x_{0},v_{0}):=(0,x,v),$
(A.2) $\begin{split}&(t_{k+1},x_{k+1},v_{k+1})\\ :=&\begin{cases}(\infty,x_{k},v_{k}),\ &(x_{k},v_{k})\in\Omega_{stop},\\ (t_{k}+t_{\partial}(x_{k},v_{k}),x_{k}-v_{k}t_{\partial}(x_{k},v_{k}),R(x_{k}-v_{k}t_{\partial}(x_{k},v_{k}),v_{k})),\ &(x_{k},v_{k})\notin\Omega_{stop},\end{cases}\end{split}$
(A.3) $n(t,x,v):=\sup\{k\in\mathbb{N}:\ t_{k}(x,v)\leq t\}.$
@@
Next, we define the last rebound from the backward trajectory of $(x,v)$:
(A.4) $(t_{fin},x_{fin},v_{fin}):=\begin{cases}(t,x_{n},v_{n}),&n(t,x,v)<\infty,\ \,t_{n(t,x,v)+1}=\infty,\\ (t_{n},x_{n},v_{n}),&n(t,x,v)<\infty,\ \,t_{n(t,x,v)+1}<\infty,\\ \lim\limits_{n\rightarrow\infty}(t_{n},x_{n},v_{n}),&n(t,x,v)=\infty.\end{cases}$
Finally, for $0\leq s\leq t$, we define the characteristic line :